find two consecutive numbers whose product is 72
Answer by dolly(163)

Let the numbers be x and x+1
Then x(x+1) = 72
==> x^2 + x - 72 = 0
==> x^2 + 9x - 8x - 72 = 0
==> x(x+9) - 8(x+9) = 0
==> (x+9) (x - 8) = 0
==> x + 9 = 0 or x - 8 = 0
==> x = -9 or x = 8
Thus the numbers are -9, -8 or 8, 9
good luck!!!
dolly